OpenStudy (anonymous):

Perimeter of the rectangle is 100 units. find length of side CD

OpenStudy (anonymous):



OpenStudy (trojanpoem):

the perimeter is AD+BC+AB+DC and 2DC+ 2BC = 2(2z+3) + 2(3z+2) 4 z + 6 + 6z + 4 = 100 10 z + 10 = 100 10 z = 90 z= 9 DC = 2* 9 + 3 = 18 + 3 = 21 BC = 3 * 9 + 2 = 29 Review for miscalculations .

OpenStudy (anonymous):

7

OpenStudy (trojanpoem):

hmm ?